OpenStudy (anonymous):

Cars arrive at a petrol station according to Poisson process with rate 24 per hour. What is the probability that no more than two cars arrive during a 5 minute interval? Select one: a. 0.3233 b. 0.6767 c. 0.2707 Incorrect d. 0.5413 e. 0.5940

OpenStudy (anonymous):

@vishweshshrimali5

OpenStudy (vishweshshrimali5):

What do you think should be the answer ? Have you tried this yet ?

OpenStudy (vishweshshrimali5):

Sorry, I can't help you in stats. Never was good at stats/

OpenStudy (anonymous):

i put c, which is incorrect

OpenStudy (anonymous):

thanks

OpenStudy (dls):

For poisson distribution,refer to this formula : \[\Huge P = \frac{\lambda^x e^{- \lambda}}{x!}\] Where lambda=Mean no. of succeses in the given time period. x : Number of succeses we are interested in e : Natural base lograrithm.

OpenStudy (anonymous):

alpha=24/60=0.4cars per min lambda=alpha*time=0.4*5=2 NOt sure from here P(X<2), am i on the right track @DLS

OpenStudy (anonymous):

no more than two means zero one two

OpenStudy (anonymous):

\[P(k=0)=e^{-\lambda}\] \[P(k=1)=\lambda e^{-\lambda}\] \[P(k=2)=\frac{\lambda^2e^{-\lambda}}{2}\]

OpenStudy (anonymous):

add them up for your answer

OpenStudy (sidsiddhartha):

for P(x<2) why ure including 2

OpenStudy (anonymous):

typo there in your case \(\lambda = 2\) since 24 per hour is the same as 2 per five minutes

OpenStudy (anonymous):

no more than two

OpenStudy (anonymous):

i drank no more than two bottles of beer last night could have had none, one, or two

OpenStudy (anonymous):

AHH!! so also include 2, i think that is where i was goin wrong

OpenStudy (anonymous):

I'm getting b. Concurr?